OpenStudy (anonymous):

Evaluate 4s^2 for s = 1/4

OpenStudy (anonymous):

\[4s^2=4\left(\frac{1}{4}\right)^2=4\frac{1}{4^2}=\frac{1}{4}\]

OpenStudy (anonymous):

thank you (: